# An Information Theory of Compute-Optimal Size Scaling, Emergence, and Plateaus in Language Models

### v1

@@ 1 Introduction
To provide simple and insightful explanations of empirical phenomena, several abstract frameworks have been proposed [2, 18, 21], all based on a skill-text bipartite graph that operates at a semantic level and captures key real-world properties [30]. Arora and Goyal [2] explain emergent phenomena by assuming a compute-optimal size scaling rule (Chinchilla allometry rule) [11]. Liao et al. [18] also assume compute-optimal (Chinchilla) size scaling to explain emergence. Michaud et al. [21] assume power-law scaling and that each text piece contains only one skill, which may be very different than real-world scenarios. Moreover, inverse polynomial loss scaling is interpreted as the average behavior of emergence at different scales. These existing frameworks explain neither the Chinchilla rule nor the plateau phenomenon. These three frameworks abstract the gradient dynamics of language model training [2]; an alternate mathematical framework considers dynamics to explain the Chinchilla rule but does not consider either emergence or plateauing [6].
@@ 3.2 Scaling of excess entropy
Using Pinsker’s inequality that relates Kullback-Leibler divergence to total variational distance as $D_{KL}(P||Q)\geq\frac{1}{2}||P-Q||$, and the equivalence between total variation distance and error rate on cloze questions [2], we obtain the following lower bound on excess entropy:
@@ 3.3 Emergence
Figure 4: Accuracy of the language model sharply increases after the model size (equivalently $C$) exceeds a threshold, which is a consequence of the emergence of a GCC in a skill graph $G_{2}^{(l)}$. (a) Step increase in accuracy for a homogeneous task. (b) Skill level distribution $q(l)$ for unimodal and multimodal heterogeneous tasks. (c) Smooth emergence for unimodal heterogeneous task. (d) Plateauing phenomena as a consequence of a task requiring diverse skills according to multimodal distribution. In this subplot, we used the following values for the parameters: number of skill levels $L=100$, $S^{(l)}=10^{3}$, $\eta_{l}=\exp(7l/L)$, $\sigma_{l}=\log_{2}(l)$ for all $l\in\{1,\ldots,L\}$, $q(m)=1/6$ for all $m\in\{2,\ldots,8\}$, .
@@
The overall accuracy is therefore a weighted average of the emergence curves. To illustrate this using a numerical example, consider a skill graph $G_{2}$ with $L=100$ levels, let $q(m,l)=q(m)q(l)$ with $q(m)=1/6$ for $m\in\{2,\ldots,8\}$ and consider a binomial distribution, $\mbox{Binomial}(L,1/2)$, over the skill levels, i.e., $q(l)=\binom{L}{l}(\frac{1}{2})^{L}$ as shown in Figure 4(b). The corresponding accuracy according to (13) is shown in Figure 4(c). In general, a smooth single phase transition can be obtained by a unimodal distribution over skill levels with a sufficiently large variance.
@@ 3.4 Plateauing
According to our framework, plateauing in accuracy after encountering an emergent phenomenon (with scaling) occurs because of the greater diversity of skills (at multiple levels) required by the heterogeneous task under consideration. In particular, we observe plateauing when the skill levels required for a task follows a multimodal distribution. To illustrate this, consider a mixture of binomial distributions over the skill levels, i.e., $q(l)=\sum_{i}w_{i}\mbox{Binomial}(L,\pi_{i})$, with $(w_{i})_{i}\in(2/5,2/5,1/5)$ and $(p_{i})_{i}=(0.2,0.6,0.95)$ is shown in Figure 4(b). The corresponding accuracy according to (13) is shown in Figure 4(d). In general, a multimodal distribution over skill levels results in emergence at multiple scales and plateaus between them. Our framework yields an interesting trend associated with the plateauing of performance: plateauing indicates the possibility of one (or more) upcoming emergent phenomenon (phenomena), which one would encounter with further scaling.
@@ A.1 A brief summary of belief propagation decoding of LDPC codes under erasure
Learning concepts from texts by the peeling process described in Section 2.3 is identical to belief propagation decoding of an LDPC code when the channel noise is erasure. To see this, treat $R$ concepts as erased codeword symbols (subset of variable nodes), and $T$ text pieces as parity checks. To obtain one-to-one correspondence, we need un-erased symbols (the remaining subset of variable nodes). Therefore, we choose (arbitrarily) a channel noise parameter $\epsilon\in(0,1)$, add $\frac{1-\epsilon}{\epsilon}R$ nodes (dummy nodes) to the set of variable nodes, and treat them as un-erased symbols. Next, add edges between every pair of dummy variable node and a parity check node with probability $p=\frac{d_{t}}{R}$. Consequently, the degree distribution of the parity check nodes (text pieces) is modified, i.e., its degree distribution is binomial with parameters $R/\epsilon$ (instead of $R$) and $d_{t}/R$, but the degree distribution of variable nodes remains unchanged. Let us call the resulting parent graph $\widetilde{G}_{1}$ [^3] In this section, we omit superscript $(C)$ in $\widetilde{G}^{(C)}_{1}$ for brevity. (see Figure 5) with the following text and concept degree distributions,
@@
Figure 5: Bipartite graph $\widetilde{G}_{1}$
@@ Appendix C Calculation of $p_{l}$
$\Pr(X\geq\eta_{l})\geq\begin{cases}\left(1-\exp\left(-\binom{R}{2}D_{KL}\left(\frac{\eta_{l}}{\binom{R}{2}}||p_{rr}\right)\right)\right)&\mbox{if }\eta_{l}\leq\binom{R}{2}p_{rr}\\ \frac{1}{\sqrt{8\eta_{l}\left(1-\frac{\eta_{l}}{\binom{R}{2}}\right)}}\exp\left(-\binom{R}{2}D_{KL}\left(\frac{\eta_{l}}{\binom{R}{2}}||p_{rr}\right)\right)&\mbox{otherwise.}\end{cases}$ (31)

### v2

@@ 1 Introduction
To provide simple and insightful explanations of empirical phenomena, several abstract frameworks have been proposed [2, 18, 21], all based on a skill-text bipartite graph that operates at a semantic level and captures key real-world properties [30]. Arora and Goyal [2] explain emergent phenomena by assuming a compute-optimal size scaling rule (Chinchilla allometry rule) [11]. Liao et al. [18] also assume compute-optimal (Chinchilla) size scaling to explain emergence. Michaud et al. [21] assume power-law scaling and that each text piece contains only one skill, which may be very different than real-world scenarios. Moreover, inverse polynomial loss scaling is interpreted as the average behavior of emergence at different scales. These existing frameworks explain neither the Chinchilla rule nor the plateau phenomenon. These three frameworks abstract the gradient dynamics of language model training [2]; an alternate mathematical framework considers dynamics to explain the Chinchilla rule and loss function plateaus but does not consider emergence[6].
@@ 3.2 Scaling of excess entropy
Using Pinsker’s inequality that relates Kullback-Leibler divergence to total variational distance as $D_{KL}(P||Q)\geq\frac{1}{2}||P-Q||_{1}^{2}$, and the equivalence between total variation distance and error rate on cloze questions [2], we obtain the following lower bound on excess entropy:
@@ 3.3 Emergence
Figure 4: Accuracy of the language model sharply increases after the model size (equivalently $C$) exceeds a threshold, which is a consequence of the emergence of a GCC in a skill graph $G_{2}^{(l)}$. (a) Step increase in accuracy for a homogeneous task. (b) Skill level distribution $q(l)$ for unimodal and multimodal heterogeneous tasks. (c) Smooth emergence for unimodal heterogeneous task. (d) Plateauing phenomena as a consequence of a task requiring diverse skills according to multimodal distribution. In this subplot, we used the following values for the parameters: number of skill levels $L=100$, $S^{(l)}=10^{3}$, $\eta_{l}=\exp(7l/L)$, $\sigma_{l}=\log_{2}(l)$ for all $l\in\{1,\ldots,L\}$, $q(m)=1/6$ for all $m\in\{2,\ldots,7\}$, .
@@
The overall accuracy is therefore a weighted average of the emergence curves. To illustrate this using a numerical example, consider a skill graph $G_{2}$ with $L=100$ levels, let $q(m,l)=q(m)q(l)$ with $q(m)=1/6$ for $m\in\{2,\ldots,7\}$ and consider a binomial distribution, $\mbox{Binomial}(L,1/2)$, over the skill levels, i.e., $q(l)=\binom{L}{l}(\frac{1}{2})^{L}$ as shown in Figure 4(b). The corresponding accuracy according to (13) is shown in Figure 4(c). In general, a smooth single phase transition can be obtained by a unimodal distribution over skill levels with a sufficiently large variance.
@@ 3.4 Plateauing
According to our framework, plateauing in accuracy after encountering an emergent phenomenon (with scaling) occurs because of the greater diversity of skills (at multiple levels) required by the heterogeneous task under consideration. In particular, we observe plateauing when the skill levels required for a task follows a multimodal distribution. To illustrate this, consider a mixture of binomial distributions over the skill levels, i.e., $q(l)=\sum_{i}w_{i}\mbox{Binomial}(L,\pi_{i})$, with $(w_{i})_{i}\in(2/5,2/5,1/5)$ and $(\pi_{i})_{i}=(0.2,0.6,0.95)$ is shown in Figure 4(b). The corresponding accuracy according to (13) is shown in Figure 4(d). In general, a multimodal distribution over skill levels results in emergence at multiple scales and plateaus between them. Our framework yields an interesting trend associated with the plateauing of performance: plateauing indicates the possibility of one (or more) upcoming emergent phenomenon (phenomena), which one would encounter with further scaling.
@@ 4 Conclusion
Acknowledgment
We appreciate discussion with and valuable suggestions from Razan Baltaji, Akhil Bhimaraju, and Moulik Choraria.
This work was supported in part by National Science Foundation grant PHY-2112890 and by DARPA grant “Modeling and Measuring Scientific Creativity”.
@@ A.1 A brief summary of belief propagation decoding of LDPC codes under erasure
Figure 5: Bipartite graph $\widetilde{G}_{1}$.
@@ Appendix C Calculation of $p_{l}$
$\Pr(X\geq\eta_{l})\geq\begin{cases}1-\exp\left(-\binom{R}{2}D_{KL}\left(\frac{\eta_{l}}{\binom{R}{2}}||p_{rr}\right)\right)&\mbox{if }\eta_{l}\leq\binom{R}{2}p_{rr}\\ \frac{1}{\sqrt{8\eta_{l}\left(1-\frac{\eta_{l}}{\binom{R}{2}}\right)}}\exp\left(-\binom{R}{2}D_{KL}\left(\frac{\eta_{l}}{\binom{R}{2}}||p_{rr}\right)\right)&\mbox{otherwise.}\end{cases}$ (31)